OpenStudy (anonymous):

Troy pushes on a car for 10 seconds, during which he applies an impulse of 300 kgm/s. What force does he apply to the car? A. 0.033 N B. 30 N C. 3000 N D. 300 N

OpenStudy (anonymous):

i think it is C. im nto 100% is that correct? @douglaswinslowcooper

OpenStudy (anonymous):

F = change in momentum / time (F)(time) = impulse = (mass)(velocity) = 300 kg-m/s F = impulse/time = (300 kg-m/s)/(10s) = 30 kg-m/s^2 = 30N Not B.

OpenStudy (anonymous):

you mean not C? @douglaswinslowcooper how would you convert kg-m/s^2 into N?

OpenStudy (anonymous):

Sorry, I meant not C. B is right, as you figured out.

OpenStudy (anonymous):

1 kg-m/s^2 = 1 J they are equal

OpenStudy (anonymous):

alright, thanks!